OpenStudy (anonymous):

PLEASE HELP! Solve the given equation 9^=35

OpenStudy (anonymous):

I don't understand the question is it 9^x = 35?

OpenStudy (anonymous):

Sorry yes 9^y=35

OpenStudy (anonymous):

???????

OpenStudy (anonymous):

You can take the log of base 9 of each side, to get: \[\log_{9}9^y = \log_{9}35\] Using laws of logarithms, the left side simplifies to y (because 9^1 = 9) And you get: \[y= \log_{9}35 \]

OpenStudy (anonymous):

Thes care my answer choices? a. 1.6181 c. 36.6783 b. 3.5553 d. 0.5898

OpenStudy (anonymous):

yln9=ln35 y=ln35/ln9 y=1.618...

OpenStudy (anonymous):

Just punch in log 35 / log 9 in your calculator..